OpenStudy (anonymous):

what is the ratio of the radius of a circle to its circumference? 1/2pi (1/pi)r 1/pi

OpenStudy (anonymous):

@superdavesuper can you help?

OpenStudy (anonymous):

is it 1/2pi

OpenStudy (ajprincess):

let the radius of the circle be r. then the circumference of the circle will be 2pi*r radius:circumference r:2pi*r r/(2pi*r) =1/(2pi)

OpenStudy (ajprincess):

yup u r right:)

OpenStudy (anonymous):

Thank you this stuff is driving me crazy. :/

OpenStudy (ajprincess):

u r welcome:) R u stil havng doubts in that?

OpenStudy (anonymous):

no i thought it was that but it seems geo is much harder than alg for me for some reason. when i think i have it right i find out im totally wrong and my logic is misplaced for some reason. So i find myself second guessing myself on everything now.

OpenStudy (ajprincess):

oh. Dnt wry u will get it right always with little more practice.:)

OpenStudy (mathmale):

I'll just summarize this discussion: The circumference of a circle is C=2(Pi)(r). We want the ratio of (the radius of the circle) to (the circumference of the circle). Algebraically, this ratio is simply r/C. But C=2(Pi) r. Then the ratio becomes r ---- , which reduces to the same 1/(2Pi) as the others have gotten. 2(Pi)r

OpenStudy (anonymous):

@mathmale thanks

OpenStudy (mathmale):

My great pleasure!